Find the complete integral of the partial differential

equation pq = 2



1
Expert's answer
2021-07-19T05:43:53-0400

We have "F(p,q)=pq-2."

Let the complete integral is given by 


"z=ax+by+c"

where


"F[a,b]=ab-2=0"

Solve for "b"


"b=\dfrac{2}{a}"

The complete integral of the of the partial differential equation "pq = 2" is



"z=ax+\dfrac{2}{a}y+c"